sinx + cosx =sqrt2 (solve on the interval [0,2pi]

Answer by lwsshak3(11628)

sinx + cosx =sqrt2 (solve on the interval [0,2pi]
try x=π/4
sin(π/4)+cos(π/4)=√2/2+√2/2=2√2/2=√2
so, x=π/4